Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
March 14, 2008

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Chris Nicolich, Amy Polk, Stephen Williams
Announcements
  1. Smash Brothers Brawl is fun.
  2. New Mario-Kart comes out on April 27.
  3. Amy's birthday was last weekend.
  4. No significant announcements from Support Services.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. The wide-screen laptop sent by GovConnection (to test with the battery charger) seems to be lost in the mail. Will has asked for a tracking number, but has not received a reply.
Liaisons
  1. Training to be scheduled for MS certified desktop support for liaisons.
  2. Will was asked for evaluation input for some of the liaisons. He passed it along to Amy.
  3. Systems training - SANS (security) - Group B is next.
Networking
  1. McAfee server - new version very close to ready. Bobbie says the server is not behaving correctly. It should really be ready next week.
  2. SAN - ongoing.
  3. CJK image - Logan spoke with David Hickey - hope to meet with him next week.
  4. The Directors are going to speak with MSL regarding the Xerox space - whether or not we can relocate the servers there.
  5. Public printing - on hold. They are only investigating solutions, not deciding on them. In addition to CIRCA, we've received proposals from Oce' and Xerox. The log-on/log-off script is done. Need to test it with AD. We could test it on the printing machine.
  6. VM backups - no consensus yet on a new place to put them. VM backup through Backup Exec - ideally, need 500G of space. Will is going to look for space.
  7. Hard drives for the Power Edge 2850 - received this week. They sent 15K RPM drives - said the 10K are discontinued. Dell says the 15K can be used with the existing 10K drives. The controller will mitigate the speed. Logan has loaded them and is currently rebuilding the server.
  8. WIN XP slipstream - Cynthia is working on the latest round of updates. It requires building a new ISO file. Trying to integrate the drivers.
  9. All current updates have been approved. Must have gone okay - have not heard any complaints. New IE7, Windows Media Player and .Net3.0 updates not approved.
Software Projects Status
  1. Web 2.0 thought the Blog Policy was very friendly. Trying out themes - going for orange and blue. Logan is copying from the Harvard site - they have explicit instructions for set up.
  2. Grover - working on the ordering page - shape, form and functionality.
  3. Tracking database - design version 4.
  4. Greenstone UFDC - Mark is working on optimization. Still getting out-of-memory exceptions.
  5. Circulation Analysis - OCLC files were not in the correct format - every other character was null. Winston fixed it.
  6. Affiliation Modification Proposal - Lori said it was acceptable. Now it needs to go to Shelley and to Judy after that - to take the proposal to Bridges.
  7. Logan met with Jason Fleming, Matthew Daley and Christine Cogar about a secure wiki for the Circ Manual. They want students to be able to browse parts of it. Circ Manual folders are secure. The solution reached: do a FAQ or a Circ manual with wiki for the public. Jason also wants a wiki for his department. Lib Guides may be able to fill the space for everyone who wants wikis - Library staff and Circ manual.
  8. Blog template discussion. Authentication between Linux and AD needs to work. That is Logan's next project to set up.
Students and Department
  1. Students should be back from Spring Break next week.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page